how do you figure out how many feet in 2 1/4 miles
Answer by stanbon(48565)

how do you figure out how many feet in 2 1/4 miles
-----------
A mile is 5280 ft.
---
(2 1/4) miles = (2 1/4)*5280 ft = (9/4)*5280 = 11880 ft
============================================================
You can find measure conversion equivalents online by using Google
Search for "feet miles"
============================================================
Cheers,
Stan H.